In 2018, Maverick Manufacturing reported sales of $1177209.0, depreciation of $18,000, EBIT of $153037.0, and net income of $67207.0. If the tax rate was 40%, what is the firm's free cash flow? Dividends are paid out at 40% of earnings for the year. The firm has 25,000 common shares outstanding.

Maverick Manufacturing
December 31 Balance Sheet

2017

2018

Assets

Cash

52000

56000

Short-term Investments

24000

16000

Accounts Receivable

164000

173000

Inventory

214000

210000

Total Current Assets

454000

455000

Net Fixed Assets (PP&E)

305000

328000

Total Assets

759000

783000

Liabilities & Owners' Equity

Accounts Payable

15180

31320

Accruals

7590

15660

Notes Payable

91080

70470

Total Current Liabilities

113850

117450

Long Term Debt

258060

237992

Total Liabilities

371910

355442

Common Stock

136620

140940

Retained Earnings

250470

286618

Owners' Equity

387090

427558

Total Liabilities & Owners' Equity

759000

783000

Fill in your answers below: 2017 2018
NOPAT
NOWC
Operating Capital
FCF
Answer #1

Free Cash flow for maverick = EBIT*(1-t) + Depreciation - Change in NWC - Capex

EBIT * (1-t) = 153037*(1-40%) = $ 91,822.2

Depreciation = 18,000

Change in NWC = NWC in 2018 - NWC in 2017 =(CA - CL) in 2018 - (CA- CL) in 2017 = (455000- 117450) - (454000- 113850) = -2600

Capex = Net fixed assets in 2018 - Net fixed assets in 2017 + Depreciation = (328000 - 305000 + 18,000) = 41,000

Now Free Cash flow = 91,822.2 + 18,000 - (-2600) - 41,000 = $ 71,422.2